On June 28, 2019, Gundrum Inc. decided to dispose of its widget-making machine and found a buyer willing to pay $15,000 for the machine. At January 1, 2019, the cost and accumulated depreciation of the widget-making machine was $1,000,000 and $900,000, respectively. The widget machine is depreciated over 20 years with no salvage value. Make all of the appropriate journal entries.
Answer #1
Date Account title Debit credit
June 28,2019 Depreciation expense 24520.55
Accumulated depreciation-widget making machine 24520.55
June 28,2019 Cash 15000
Accumulated depreciation-widget making machine(900000+24520.55) 924520.55
loss on sale of widget making machinery 60479.45
Widget making machine 1,000,000

**

Depreciation =[cost-salvage value]/useful life

        =[1,000,000-0]/20

        = 1,000,000/20

        = 50000

Depreciation for 6months(1Jan -28June or 179 days )= 50000*6/12= 25000   or 50000*179/365= 24520.55

b)Book value =cost-accumulated depreciation

                     = 1000000-924520.55

                    = 75479.45

Loss on sale =15000-75479.45 = 60479.45